Ask your own question, for FREE!
Mathematics 13 Online
OpenStudy (anonymous):

A bag contains 20 checkers -- 10 red and 10 black. Find the probability of selecting 2 black checkers when they are chosen at randon with replacement. Determine whether the event is dependent or independent.

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

what's the chance 1st pull?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

small rant: an event cannot be dependent or independent dependent or independent is relation between two events

OpenStudy (fibonaccichick666):

good, now second pull?

OpenStudy (anonymous):

1/4 then and independent since it doesn't matter since its being replaced 1/2*1/2=1/4

OpenStudy (fibonaccichick666):

yup, you got it

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
3 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!